# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v2

@@ 3.2.1 RU-MLMC
Recall that both the loss function and its gradient can be represented as the summation of the expected value of the random variable sequence, and this topic has been well studied. We introduce a non-negative integer-valued random variable $L$ that is independent of the two random variable sequences, with its probability mass function $\mathbb{P}(L=\ell)=w_{\ell}$. By the law of total expectation, we have
@@
$V_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\rho_{\phi,L}}$ (22)
leads to an unbiased Monte Carlo estimator of the gradient $\nabla_{\phi}\mathcal{L}(\phi)$.
@@
$U_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\psi_{\phi,L}}$
for the query of the estimation of the loss function $\mathcal{L}(\phi)$. This method is known as the randomized unbiased multilevel Monte Carlo (RU-MLMC) method [51]. In this study, we take $L$ as a geometric distribution $\mathrm{Ge}(p)$ with $\omega_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. To ensure a finite variance and finite expected computational cost for $V_{\mathrm{RU}},U_{\mathrm{RU}}$, it is required that $\alpha\in(1,\min\left(r_{1},r_{2}\right))$, where $r_{1}$ and $r_{2}$ are from Theorem 3.1 and Theorem 3.2, respectively. This is further examined with the following theorem. Since the gradient $V_{\mathrm{RU}}$ is a vector, the notation $\mathrm{Var}[V_{\mathrm{RU}}]$ used in the following applies to the variance of each component of $V_{\mathrm{RU}}$ for simplicity. Denote the cost of RU-MLMC as $\mathrm{Cost}_{\mathrm{RU}}$.
@@
Bounding $\mathrm{Var}[V_{\mathrm{RU}}]$ is just a replica by using Theorem 3.2.
@@ 3.2.2 GRR-MLMC
The Russian Roulette (RR) estimator [43] is also employed to estimate the sum of an infinite series, wherein the evaluation of any term in the series only demands a finite amount of computation. This estimator relies on randomized truncation and assigns a higher weight to each term to accommodate the possibility of not computing them. The query of the gradient of the RR-MLMC estimator is given by
@@
Motivated by [42], to trade lower variance with higher cost based on the conventional RR-MLMC, one way is to ensure that the first ${\underline{m}}$ terms of the infinite series are always computed, where this ${\underline{m}}$ is called the base level. In this case, the random index $L$ is set to have a lower bound ${\underline{m}}$, i.e., $\mathbb{P}(L\geq{\underline{m}})=1$, implying $p_{j}=1$ for all $j\leq{\underline{m}}$. We call this modified estimator the generalized Russian roulette (GRR) estimator. When ${\underline{m}}=0$, this degenerates to the conventional RR-MLMC. The associated query for the GRR-MLMC estimator of the gradient is
@@
As we can predict, a larger $\underline{m}$ leads to lower variances, but it increases the costs. On the other hand, for a fixed $\underline{m}$, a larger $\alpha$ leads to a lower cost, but it raises the variances.
@@ 3.3 Truncated MLMC for APT
When treating $\nabla_{\phi}\mathcal{L}(\phi)$ as the summation of the infinite random variable sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, we notice that $\Delta\rho_{\phi,0}=\rho_{\phi,0}$ is just the nested APT estimator with a bias of order $\mathcal{O}(1/M_{0})$. For $\Delta\rho_{\phi,\ell}$ with $\ell\geq 1$, it actually contributes to the reduction of bias while simultaneously increasing the variance and this additional variance grows with $\ell$. An empirical demonstration of the additional variance for $\Delta\rho_{\phi,\ell}$ is shown in Figure 1.
@@
We observe that the truncated RU-MLMC method enjoys a more stable loss compared to the original one, indicating that truncation helps reducing variance. Therefore, in order to mitigate the variance of the RU-MLMC, we propose to truncate the distribution of $L$ by setting the largest value of $L$ as $\overline{m}$. The truncation leads to a bias of order $\mathcal{O}(1/M_{{\overline{m}}})$. Following [32], we extend the truncated idea to other unbiased MLMC methods.
@@ 3.3.1 TGRR-MLMC
R-MLMC. Let ${\overline{m}}\geq{\underline{m}}$ be a truncated level, and let $L$ be a nonnegative integer-value random index taking values in $\{{\underline{m}},\dots,{\overline{m}}\}$. The query for the estimation of the gradient of TGRR-MLMC is then defined as
@@
$=\mathbb{E}[\rho_{\phi,M_{{\underline{m}}}}]+\sum_{j={\underline{m}}+1}^{{\overline{m}}}\frac{\mathbb{E}[\Delta\rho_{\phi,j}]}{p_{j}}\sum_{\ell\geq j}^{{\overline{m}}}\mathbb{P}(\tilde{L}=\ell)$
@@
We should note that if ${\overline{m}}={\underline{m}}=m$, TGRR-MLMC degenerates to nested APT with the inner sample size $M_{m}=M_{0}2^{m}$. For any truncated method with truncated level $\overline{m}$, they all share the bias of order $\mathcal{O}(2^{-\overline{m}})$.
@@
where $w_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. We denote this distribution by $\mathrm{Ge}(p,{\underline{m}},{\overline{m}})$. The whole procedure of TGRR-MLMC method is summarized in Algorithm 1.
Input: Prior $p(\theta)$, implicit simulator model $p(x|\theta)$, conditional density estimator $q_{F(x,\phi)}(\theta)$, $\alpha\in(1,\min(r_{1},r_{2}))$, $M_{0}$, $M_{\ell}=M_{0}2^{\ell}$, truncated level ${\overline{m}}$, base level ${\underline{m}}$, total round $R$, optimizer $\mathrm{opt}(\cdot,\cdot)$, batch size $B$, outer sample size $N$
@@ 4 Convergence results of SGD
In this section, we analyze the convergence of the methods proposed in this paper with some existing techniques. For simplicity, we only study the widely used SGD with constant step size $\gamma$ in this paper. Our result should extend to border cases. The parameter update procedure of SGD is presented as
@@
where $V_{\mathrm{TGRR}}^{(i)}$ are iid copies of $V_{\mathrm{TGRR}}$. The gradient estimators for other MLMC methods follows a similar form. The gradient estimator for nested APT is stated previously in (16). Since some of them are biased estimators for $\nabla_{\phi}\mathcal{L}(\phi)$, we take the following decomposition
@@
where $b(\phi)=\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]-\nabla_{\phi}\mathcal{L}(\phi)$ and $\eta(\phi)=\nabla_{\phi}\hat{\mathcal{L}}(\phi)-\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]$ denote the bias and the noise of gradient estimator $\nabla_{\phi}\hat{\mathcal{L}}(\phi)$, respectively.
@@
If the SGD algorithm (28) has a step size $\gamma\leq\min(1/K,1/\mu)$, for a fixed $T$ iterations, the upper bound of the optimal gap (30) is given as
@@
This is a special case of Theorem 6 in [1] by taking $m=M=0$ there.
Since $G_{0}$ depends solely on the initialization of $\phi$, our focus lies on the last two terms of (33), where $U_{b}$ and $U_{\eta}$ have the same impact on the optimal gap $G_{T}$. This implies that unbiased methods with large variance may ultimately achieve the same result as those biased ones at the same iteration step $T$. Hence in cases where variance dominates bias, methods with smaller variance are preferable. On the other hand, when a sufficiently large iteration step $T$ is given and proper learning rate $\gamma$ is chosen, the upper bound of the optimal gap can be smaller than any given $\epsilon>0$ except for the bias, which is examined in the second part of theorem 4.2. In this case, unbiased methods are always favorable, as they can converge to an arbitrarily small neighborhood of 0, while biased methods can only converge to a neighborhood of the $U_{b}/(2\mu)$. Moreover, for biased methods with $U_{b}\gg U_{\eta}$, the optimal gap for $t_{1}\gg t_{2}$ tends to be $G_{t_{1}}\approx G_{t_{2}}$, indicating that increasing the iteration step can be futile in attempting to reduce the optimal gap.
@@ 5 Numerical experiments
In this section, numerical experiments compare the efficiency of two unbiased MLMC methods and a biased MLMC method for three models. To handle high-dimensional datasets, it is common to use informative low-dimensional summary statistics of the datasets instead. However, this introduces additional bias, as detailed in [14]. In this paper, we do not look into the effects of using summary statistics for SNPE methods. We refer to Fearnhead and Prangle [14] for a semi-automatic method of constructing summary statistics in the context of ABC.
@@
The intermediate variables $p$ follow a single crescent-shaped distribution, which is then shifted and rotated around the origin based on the parameter values of $\theta$. The absolute value $|\theta_{1}+\theta_{2}|$ contributes to the emergence of a second crescent in the posterior distribution. We choose a uniform prior over the square $[-1,1]^{2}$ to perform the inference. The observation for this task is $x_{o}=(0,0)$ following [24], the ground truth parameters are $\theta^{*}=(0.2475,0.2475).$
Lotka-Volterra model. This model describes the continuous time evolution of a population of predators interacting with a population of prey using a stochastic Markov jump process. The model describes that the birth of a predator at a rate $\exp(\theta_{1})XY$, resulting in an increase of $X$ by one. The death of a predator at a rate proportional to $\exp(\theta_{2})X$, leading to a decrease of $X$ by one. The birth of a prey at a rate proportional to $\exp(\theta_{3})Y$, resulting in an increase of $Y$ by one. The consumption of a prey by a predator at a rate proportional to $\exp(\theta_{4})XY$, leading to a decrease of $Y$ by one. Following the experimental details outlined in [48], we initialize the predator and prey populations as $X=50$ and $Y=100$, respectively. We conduct simulations of the Lotka-Volterra model using Gillespie’s algorithm [20] throughout 30 time units. We recorded the populations at intervals of 0.2 time units, resulting in time series data sets, each consisting of 151 values. The resulting summary statistics $S(x)$ are represented as a 9-dimensional vector, which includes the following time series features: the logarithm of the mean of each time series, the logarithm of the variance of each time series, the auto-correlation coefficient of each time series at lags of 0.2 and 0.4-time units, and the cross-correlation coefficient between the two time series. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, and we generate the ground truth posterior with SMC-ABC [4], which is more costly than methods in this paper. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, the ground truth parameters are
@@
and the corresponding standard deviation under the ground truth parameters $\theta^{*}$ (based on 10,000 simulations) are
[…]
M/G/1 queue model. The M/G/1 queue model [55] describes a single server’s processing of a queue of continuously arriving jobs. Define $I$ as the total number of jobs that needs to be processed, and denote by $s_{i}$ the processing time required for job $i$. Let $v_{i}$ be the job’s arrival time in the queue, and $d_{i}$ be the job’s departure time from the queue. They satisfy the following conditions
@@
In our experiments, we set $I=50$ jobs and the summary statistics $S(x)$ has been selected as the logarithm of 0th, 25th, 50th, 75th and 100th percentiles of the set of inter-departure times. The prior distribution of the parameters is given by
@@
[…]
@@
[…]
Our numerical experiments are performed on a computer equipped with a single GeForce RTX 2080s GPU and an i9-9900K CPU. The training and inference processes of the model are primarily implemented using the Pytorch package in Python.
@@
Figure 2: Approximated posterior for RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-moon, from left to right: available ground truth, RU-MLMC, GRR-MLMC and TGRR-MLMC. B. Lotka-Volterra, from left to right: ground truth simulated with SMC-ABC [4], RU-MLMC, GRR-MLMC and TGRR-MLMC. C. M/G/1 queue model, the setting is the same with Lotka-Volterra. We have plotted the scatter plot for each of the 2D sub-spaces. The histogram for each $\theta_{i},i=1,\dots,4$ is plotted on the diagonal with the ground truth parameter marked with dotted lines.
@@
In the training process, we simulate $N=1,000$ samples in each round, and with $R=20$ rounds in total. In each round, we randomly pick 5% of the newly generated samples $\theta$ and their corresponding $x$ values as validation data. We follow the early stop criterion proposed by [48], which terminates the training if the loss value on the validation data does not decrease after 20 epochs in a single round. For the optimizer, we use Adam [35] with a batch size of 100, a learning rate of $1\times 10^{-4}$, and a weight decay of $1\times 10^{-4}$. The detailed track of the computational time of various methods can be found in Table 3.
@@
In this paper, we employ neural spline flows (NSFs) [13] as the conditional density estimator, which consists of 8 layers. Each layer is constructed using two residual blocks with 50 units and ReLU activation function. With 10 bins in each monotonic piecewise rational-quadratic transform the tail bound is set to 20. For a detailed introduction to the structure of NSFs, we refer the reader to Appendix B.
@@
We compare the performance of RU-MLMC, GRR-MLMC, and TGRR-MLMC on the Two-moon, Lotka-Volterra, M/G/1 models. To assess the similarity between the approximate posterior distribution $q_{F(x_{o},\phi)}(\theta)$ and the true posterior distribution $p(\theta|x_{o})$ given the observed data, we employ maximum mean discrepancy (MMD) [24, 25, 30, 48] and classifier two-sample tests (C2ST) [11, 27, 39] as discriminant criteria. Furthermore, we use the logarithmic median distance (LMD) to measure the distance between $x_{o}$ and $x$ drawn from $p(x|\theta)$, where $\theta$ is sampled from $q_{F(x_{o},\phi)}(\theta)$. The negative log-density (NLOG) value of the ground truth parameter $\theta^{*}$ of the generated observation $x_{o}$ under the conditional density structure, i.e., $-\log q_{F(x_{o},\phi)}(\theta^{*})$, is another useful metric. A lower value indicates that, in the parameter space, the posterior density estimator has a higher probability density at the true parameter, which can be used to evaluate the model’s accuracy in estimating the true parameter given the observed data [13, 24, 30, 47, 48].
@@
For the setting of MLMC methods, we take $M_{0}=8$ and $\overline{m}=4$ for TGRR-MLMC. The base level for GRR and TGRR is ${\underline{m}}=2$. Since the proposal distributions of each round are distinct, the hyperparameter $r_{2}$ in Theorem 3.2 differs in each round. To determine the value of $r_{2}$ for each round, except for the first two, one may take the value from the previous round. To address this issue, we have conducted 50 training processes and sequentially performed linear regression, a universal value is then selected as $r_{2}=1.8$.
@@
We now discuss the choice of the hyperparameter $\alpha$ for the geometric distribution of the level $L$. Unlike previous work where $\alpha$ is chosen to minimize the average cost $\mathrm{Cost}_{\mathrm{RU}}$ [21, 22, 29], based on experiment results in Figure 1, in addition to average cost, variance is another crucial factor that demands our attention. These two indicators collectively hold great significance in our evaluation and optimization processes. We now try to minimize the asymptotic inefficiency [34] defined as $H_{\mathrm{RU}}:=\mathrm{Var}[V^{\mathrm{RU}}]\times\mathrm{Cost}_{\mathrm{RU}}$. Utilizing Theorem 3.3, we find an upper bound for the asymptotic inefficiency
@@
for some constant $B>0$. As a result, the optimal $\alpha$ minimizing the upper bound is $\alpha_{\mathrm{RU}}^{*}=(r_{2}+1)/2\in(1,r_{2})$. When ${\underline{m}}=0$ in GRR-MLMC, which degenerates to RR-MLMC, we reach the same conclusion for $\alpha^{*}_{\mathrm{RR}}$. From this point of view, we can see that RR-MLMC actually trades lower variance for higher average cost compared with RU-MLMC. We apply the same procedure for other MLMC methods, and the corresponding optimal $\alpha$ are presented in Table 1. For an ablation study on the choices of $\alpha$, we refer the reader to Appendix D.
Table 1: The value of alpha minimizing the upper bound of asymptotic inefficiency

### v3

@@ 3.2.1 RU-MLMC
Recall that both the loss function and its gradient can be represented as a sum of expectations over a sequence of random variables; this topic has been well studied. We introduce a non-negative integer-valued random variable $L$ that is independent of the two random variable sequences, with its probability mass function $\mathbb{P}(L=\ell)=w_{\ell}$. By the law of total expectation, we have
@@
$V_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\rho_{\phi,L}},$ (22)
leads to an unbiased MLMC estimator of the gradient $\nabla_{\phi}\mathcal{L}(\phi)$.
@@
$U_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\psi_{\phi,L}},$
for the estimation of the loss function $\mathcal{L}(\phi)$. This method is known as the randomized unbiased multilevel Monte Carlo (RU-MLMC) method [51]. In this study, we take $L$ as a geometric distribution $\mathrm{Ge}(p)$ with $\omega_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$ [22]. To ensure a finite variance and finite expected computational cost for $V_{\mathrm{RU}},U_{\mathrm{RU}}$, it is required that $\alpha\in(1,\min\left(r_{1},r_{2}\right))$, where $r_{1}$ and $r_{2}$ are from Theorem 3.1 and Theorem 3.2, respectively. This is further examined with the following theorem. Since the gradient $V_{\mathrm{RU}}$ is a vector, the notation $\mathrm{Var}[V_{\mathrm{RU}}]$ used in the following applies to the variance of each component of $V_{\mathrm{RU}}$ for simplicity. Denote the cost of RU-MLMC by $\mathrm{Cost}_{\mathrm{RU}}$.
@@
Bounding $\mathrm{Var}[V_{\mathrm{RU}}]$ follows analogously from Theorem 3.2.
@@ 3.2.2 GRR-MLMC
RU-MLMC methods can at times suffer from excessive variance, which in turn degrades convergence behavior [22, 29]. To address this, we employ alternative schemes that reduce variance while preserving unbiasedness. The Russian Roulette (RR) estimator [43] is also employed to estimate the sum of an infinite series, wherein the evaluation of any term in the series only demands a finite amount of computation. This estimator relies on randomized truncation and assigns a higher weight to each term to accommodate the possibility of not computing them. The gradient estimator of RR-MLMC is given by
@@
Motivated by [42], to trade lower variance for higher cost based on the conventional RR-MLMC, one way is to ensure that the first $\underline{m}$} terms of the infinite series are always computed, where this $\underline{m}$} is called the base level. In this case, the random index $L$ is set to have a lower bound $\underline{m}$}, i.e., $\mathbb{P}(L\geq{\underline{m}})=1$, implying $p_{j}=1$ for all $j\leq{\underline{m}}$. We call this modified estimator the generalized Russian roulette (GRR) estimator. When ${\underline{m}}=0$, this degenerates to the conventional RR-MLMC. The associated query for the GRR-MLMC estimator of the gradient is
@@
As expected, a larger $\underline{m}$ leads to lower variances, but it increases the costs. On the other hand, for a fixed $\underline{m}$, a larger $\alpha$ leads to a lower cost, but it raises the variances.
@@ 3.3 Truncated MLMC for APT
In practice, a truncated variant of the unbiased method can be preferable, which introduces a small, controlled bias that yields lower variance while also bounding computational cost and peak GPU memory, making the estimators more stable under fixed compute budgets. When treating $\nabla_{\phi}\mathcal{L}(\phi)$ as the summation of the infinite random variable sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, we notice that $\Delta\rho_{\phi,0}=\rho_{\phi,M_{0}}$ is just the nested APT estimator with a bias of order $\mathcal{O}(1/M_{0})$. For $\Delta\rho_{\phi,\ell}$ with $\ell\geq 1$, it actually contributes to the reduction of bias while simultaneously increasing the variance, and this additional variance grows with $\ell$. An empirical demonstration of the additional variance for $\Delta\rho_{\phi,\ell}$ is shown in Figure 1.
@@
We observe that the truncated RU-MLMC method enjoys a more stable loss than the original one, indicating that truncation helps reduce variance. Therefore, to mitigate the variance of the RU-MLMC, we propose to truncate the distribution of $L$ by setting the largest value of $L$ as $\overline{m}$. The truncation leads to a bias of order $\mathcal{O}(1/M_{{\overline{m}}})$. Following [32], we extend the truncated idea to other unbiased MLMC methods.
@@ 3.3.1 TGRR-MLMC
R-MLMC. Let ${\overline{m}}\geq{\underline{m}}$ be a truncated level, and let $L$ be a random index taking values in $\{{\underline{m}},\dots,{\overline{m}}\}$. The query for the estimation of the gradient of TGRR-MLMC is then defined as
@@
[…]
@@
We should note that if ${\overline{m}}={\underline{m}}=m$, TGRR-MLMC degenerates to nested APT with the inner sample size $M_{m}=M_{0}2^{m}$. All truncated methods with truncated level $\overline{m}$ share the bias of order $\mathcal{O}(2^{-\overline{m}})$.
@@
where $w_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. We denote this distribution by $\mathrm{Ge}(p,{\underline{m}},{\overline{m}})$. The whole procedure of the TGRR-MLMC method is summarized in Algorithm 1.
Input: Prior $p(\theta)$, implicit simulator model $p(x|\theta)$, conditional density estimator $q_{F(x,\phi)}(\theta)$, $\alpha\in(1,\min(r_{1},r_{2}))$, $_{0}, M=M$$_{\ell}=M_{0}2^{\ell}$, truncated level $\overline{m}$}, base level $\underline{m}$}, total round $R$, optimizer $\mathrm{opt}(\cdot,\cdot)$, batch size $B$, outer sample size NN
@@ 31  Update proposal $\tilde{p}_{k+1}(\theta)\leftarrow q_{F(x_{o},\phi_{t})}(\theta)$
In this section, we develop three MLMC estimators for the nested APT objective: two unbiased (RU-MLMC, GRR-MLMC) and one biased (TGRR-MLMC). RU-MLMC has the lowest overhead but higher variance; GRR-MLMC reduces variance by deterministically computing the first $\underline{m}$ levels before randomized truncation; TGRR-MLMC deterministically truncates the tail to control runtime and GPU memory, introducing a small, controllable bias, retaining a favorable variance–cost behavior. These methods expose clear bias–variance–cost trade-offs, making each preferable under different compute regimes.
@@ 4 Convergence results of SGD
In this section, we analyze the convergence of the methods proposed in this paper with some existing techniques. For simplicity, we only study the widely used SGD with constant step size $\gamma$ in this paper. Our results should extend to broader cases. The parameter update procedure of SGD is presented as
@@
where $V_{\mathrm{TGRR}}^{(i)}$ are iid copies of $V_{\mathrm{TGRR}}$. The gradient estimators for other MLMC methods follow a similar form. The gradient estimator for nested APT is stated previously in (16). Since some of them are biased estimators for $\nabla_{\phi}\mathcal{L}(\phi)$, we take the following decomposition
@@
where $b(\phi)=\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]-\nabla_{\phi}\mathcal{L}(\phi)$ and $\eta(\phi)=\nabla_{\phi}\hat{\mathcal{L}}(\phi)-\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]$ denote the bias and the noise of gradient estimator by $\nabla_{\phi}\hat{\mathcal{L}}(\phi)$, respectively.
@@
If the SGD algorithm (28) has a step size $\gamma\leq\min(1/K,1/\mu)$, for a fixed $T$ iterations, the upper bound of the optimal gap (30) is given by
@@
This is a special case of Theorem 6 in [1] by taking $m=M=0$ there. ∎
Since $G_{0}$ depends solely on the initialization of $\phi$, our focus lies on the last two terms of (33), where $U_{b}$ and $U_{\eta}$ have the same impact on the optimal gap $G_{T}$. This implies that unbiased methods with large variance may ultimately achieve the same result as those biased ones at the same iteration step $T$. Hence, in cases where variance dominates bias, methods with smaller variance are preferable. On the other hand, when a sufficiently large iteration step $T$ is given and a proper learning rate $\gamma$ is chosen, the upper bound of the optimal gap can be smaller than any given $\epsilon>0$ except for the bias, which is examined in the second part of theorem 4.2. In this case, unbiased methods are more favorable, as they can converge to an arbitrarily small neighborhood of 0, while biased methods can only converge to a neighborhood of the $U_{b}/(2\mu)$. Moreover, for biased methods with $U_{b}\gg U_{\eta}$, the optimal gap for $t_{1}\gg t_{2}$ tends to be $G_{t_{1}}\approx G_{t_{2}}$, indicating that increasing the iteration step can be futile in attempting to reduce the optimal gap.
@@ 5 Numerical experiments
In this section, numerical experiments compare the performance of two unbiased MLMC methods and a biased MLMC method for three models. To handle high-dimensional datasets, it is common to use informative low-dimensional summary statistics of the datasets instead. However, this introduces additional bias, as detailed in [14]. In this paper, we do not look into the effects of using summary statistics for SNPE methods. We refer to Fearnhead and Prangle [14] for a semi-automatic method of constructing summary statistics in the context of ABC.
@@
The intermediate variables $p$ follow a single crescent-shaped distribution, which is then shifted and rotated around the origin based on the parameter values of $\theta$. The absolute value $|\theta_{1}+\theta_{2}|$ contributes to the emergence of a second crescent in the posterior distribution. We choose a uniform prior over the square $1,1]^{2}$ to perform the inference. The observation for this task is $x_{o}=(0,0)$ following [24], the ground truth parameters are $\theta^{*}=(0.2475,0.2475).$
Lotka-Volterra model. This model describes the continuous time evolution of a population of predators interacting with a population of prey using a stochastic Markov jump process. The model describes that the birth of a predator at a rate $\exp(\theta_{1})XY$, resulting in an increase of $X$ by one. The death of a predator at a rate proportional to $\exp(\theta_{2})X$, leading to a decrease of $X$ by one. The birth of a prey at a rate proportional to $\exp(\theta_{3})Y$, resulting in an increase of $Y$ by one. The consumption of a prey by a predator at a rate proportional to $\exp(\theta_{4})XY$, leading to a decrease of $Y$ by one. Following the experimental details outlined in [48], we initialize the predator and prey populations as $X=50$ and $Y=100$, respectively. We conduct simulations of the Lotka-Volterra model using Gillespie’s algorithm [20] throughout 30 time units. We recorded the populations at intervals of 0.2 time units, resulting in time series data sets, each consisting of 151 values. The resulting summary statistics $S(x)$ are represented as a 9-dimensional vector, which includes the following time series features: the logarithm of the mean of each time series, the logarithm of the variance of each time series, the auto-correlation coefficient of each time series at lags of 0.2 and 0.4 time units, and the cross-correlation coefficient between the two time series. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, and we generate the ground truth posterior with SMC-ABC [4], which is more costly than the methods in this paper. The ground truth parameters are
@@
and the corresponding standard deviation under the ground truth parameters $\theta^{*}$ (based on 10,000 simulations) is
s=(0.3294, 0.5483, 0.6285, 0.9639, 0.0091, 0.0222, 0.0107, 0.0224, 0.1823).s=(0.3294,\ 0.5483,\ 0.6285,\ 0.9639,\ 0.0091,\ 0.0222,\ 0.0107,\ 0.0224,\ 0.1823).
M/G/1 queue model. The M/G/1 queue model [55] describes a single server’s processing of a queue of continuously arriving jobs. Define $I$ as the total number of jobs that need to be processed, and denote by $s_{i}$ the processing time required for job $i$. Let $v_{i}$ be the job’s arrival time in the queue, and $d_{i}$ be the job’s departure time from the queue. They satisfy the following conditions
@@
In our experiments, we set $I=50$ jobs and the summary statistics $S(x)$ have been selected as the logarithm of 0th, 25th, 50th, 75th, and 100th percentiles of the set of inter-departure times. The prior distribution of the parameters is given by
@@
and our experiments choose ground truth parameters as $\theta^{*}=(1,\ 4,\ 0.2).$ The observed summary statistic $S(x_{o})$ simulated from the model with the ground truth parameters $\theta^{*}$ is
$S(x_{o})=(0.0929,\ 0.8333,\ 1.4484,\ 1.9773,\ 3.1510),$
@@
s=(0.1049, 0.1336, 0.1006, 0.1893, 0.2918).s=(0.1049,\ 0.1336,\ 0.1006,\ 0.1893,\ 0.2918).
Our numerical experiments are performed on a computer equipped with a single GeForce RTX 2080 Super GPU and an i9-9900K CPU. The training and inference processes of the model are primarily implemented using the PyTorch package in Python.
@@
Figure 2: Approximated posterior for RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-Moon, from left to right: available ground truth, RU-MLMC, GRR-MLMC, and TGRR-MLMC. B. Lotka-Volterra, from left to right: ground truth simulated with SMC-ABC [4], RU-MLMC, GRR-MLMC, and TGRR-MLMC. C. M/G/1 queue model, the setting is the same as Lotka-Volterra. We show a scatter plot for each of the 2D subspaces. The histogram for each $\theta_{i},i=1,\dots,4$ is plotted on the diagonal with the ground truth parameter marked with dotted lines.
@@
In the training process, we simulate $N=1,000$ samples in each round, and with $R=20$ rounds in total. In each round, we randomly pick 5% of the newly generated samples $\theta$ and their corresponding $x$ values as validation data. We follow the early stopping criterion proposed by [48], which terminates the training if the loss value on the validation data does not decrease after 20 epochs in a single round. For the optimizer, we use Adam [35] with a batch size of 100, a learning rate of $1\times 10^{-4}$, and a weight decay of $1\times 10^{-4}$. The detailed track of the computational time of various methods can be found in Table 3.
@@
In this paper, we employ neural spline flows (NSFs) [13] as the conditional density estimator, which consists of 8 layers. Each layer is constructed using two residual blocks with 50 units and ReLU activation function. With 10 bins in each monotonic piecewise rational-quadratic transform, the tail bound is set to 20. For a detailed introduction to the structure of NSFs, we refer the reader to Appendix B.
@@
We compare the performance of RU-MLMC, GRR-MLMC, and TGRR-MLMC on the Two-Moon, Lotka-Volterra, M/G/1 models. To assess the similarity between the approximate posterior distribution $q_{F(x_{o},\phi)}(\theta)$ and the true posterior distribution $p(\theta|x_{o})$ given the observed data, we employ maximum mean discrepancy (MMD) [24, 25, 30, 48] and classifier two-sample tests (C2ST) [11, 27, 39] as discriminant criteria. Furthermore, we use the logarithmic median distance (LMD) to measure the distance between $x_{o}$ and $x$ drawn from $p(x|\theta)$, where $\theta$ is sampled from $q_{F(x_{o},\phi)}(\theta)$. The negative log-density (NLOG) value of the ground truth parameter $\theta^{*}$ of the generated observation $x_{o}$ under the conditional density structure, i.e., $\log q_{F(x_{o},\phi)}(\theta^{*})$, is another useful metric. A lower value of NLOG indicates that, in the parameter space, the posterior density estimator has a higher probability density at the true parameter, which can be used to evaluate the model’s accuracy in estimating the true parameter given the observed data [13, 24, 30, 47, 48].
@@
For the setting of MLMC methods, we take $M_{0}=8$ and $\overline{m}=4$ for TGRR-MLMC. The base level for GRR and TGRR is ${\underline{m}}=2$. Since the proposal distributions of each round are distinct, the hyperparameter $r_{2}$ in Theorem 3.2 differs in each round. To determine the value of $r_{2}$ for each round, except for the first two, one may take the value from the previous round. To address this issue, we have conducted 50 training processes and sequentially performed linear regression. From the results, we select a global value of $r_{2}=1.8$.
@@
We now discuss the choice of the hyperparameter $\alpha$ for the geometric distribution of the level $L$. Unlike previous work where $\alpha$ is chosen to minimize the average cost $\mathrm{Cost}_{\mathrm{RU}}$ [21, 22, 29], based on experiment results in Figure 1, in addition to average cost, variance is another crucial factor that demands our attention. Both metrics matter in our evaluation and optimization processes. We now try to minimize the asymptotic inefficiency [34] defined as $H_{\mathrm{RU}}:=\mathrm{Var}[V_{\mathrm{RU}}]\times\mathrm{Cost}_{\mathrm{RU}}$. Utilizing Theorem 3.3, we find an upper bound for the asymptotic inefficiency
@@
for some constant $B>0$. As a result, the optimal $\alpha$ minimizing the upper bound is $\alpha_{\mathrm{RU}}^{*}=(r_{2}+1)/2\in(1,r_{2})$. When ${\underline{m}}=0$ in GRR-MLMC, which degenerates to RR-MLMC, we reach the same conclusion for $\alpha^{*}_{\mathrm{RR}}$. From this point of view, we can see that RR-MLMC actually trades lower variance for higher average cost compared with RU-MLMC. We apply the same procedure for other MLMC methods, and the corresponding optimal $\alpha$ are presented in Table 1. For an ablation study on the choices of $\alpha$, we refer the readers to Appendix D.
Table 1: The value of $\alpha$ minimizing the upper bound of asymptotic inefficiency